Problem 6 Solve the following IVP: a) y – 4y +20y=0 ; y0)=2 , y(0)=0, y(0) = 6 b) y +4y=0; y(0)=2 , y(0) = 3
Answer #1

* IVS: a) y ... 49 tarmy = 0 : Y()- 2, Y (0)= (, Y0)=-6. wolite above equation as:- y=> 3, Y=;?, Y=r. Su. 73-47?+208C,= 2, C2 = -2 Purt this value in equation . 9(+) = el* ( 2 C084+ - 2 Sin4+) = 2.e2+(Cos 4+ -Sin 4t) | 9(+) = 2.elt (cogut-Si